Gases in Chemical Reactions:

Part A
In the reaction
2Ag2O(s)→4Ag(s)+O2(g)
4.58 L of O2 was formed at P=745mmHg and T=308K. How many grams of Ag2O decomposed?
Express the mass in grams to three significant figures.

Part B
In the previous reaction, what mass of Ag2O(s) (in grams) is required to form 388.00 mL of oxygen gas at P=734mmHg and T=25∘C?
Express the mass in grams to three significant figures.
1
Expert's answer
2017-03-13T11:10:06-0400
A)
PV = mRT /M; pV=nRT; n=PV/RT

n = (99325.1645Pa*0.00458m³)/8.314J/MolK*308K = 0.1776Mol

From equation
xMolAg₂O/2MolAg₂O=0.1776MolO₂/1MolO₂

n(Ag₂O)=0.3552Mol

m(Ag₂O)=n*M=0.3552Mol*232g/Mol=82.401g

B )
n=97858.6Pa*0.000388m³/8.314J/MolK * 298ᵒK=0.0153Mol

From equation
xMolAg₂O/2MolAg₂O=0.0153MolO₂/1MolO₂

n(Ag₂O)=0.0306Mol

m(Ag₂O)=n*M=0.0306Mol*232g/Mol=7.099g